OpenStudy (abbycross167):

Can someone please help me? I really need help... I will fan and Medal Question~ A farmer has 2 irrigation sprinklers; one covers a circle with radius r and the other covers a circle with radius 4r. Find the ratio of the smaller irrigation circle to the area of the larger irrigation circle.

jhonyy9 (jhonyy9):

do you know formula for area of a circle ?

OpenStudy (abbycross167):

A=pr^2 ?

jhonyy9 (jhonyy9):

pi*r^2 =area of a circle yes ?

OpenStudy (abbycross167):

I thought it was A=pr^2?

jhonyy9 (jhonyy9):

area of circle with radius r = pi *r^2 where pi=3,14 area of circle with radius 4r = pi *(4r)^2 where pi=3,14 ratio area of smaller to area the larger = pi*r^2 / pi*(4r)^2 = = pi*r^2 / pi * 16r^2 = 1/16 ratio area smaller to the area larger circle = 1/16

OpenStudy (abbycross167):

thank you very much!! you made this make so much more sense!

jhonyy9 (jhonyy9):

np welcome and was my pleasure and anytime when i m here online just tag me there please good luck bye bye